OpenStudy (anonymous):

Find the exact circumference, in terms of pi, of a circle with the given radius. 7 feet

myininaya (myininaya):

C=2*pi*r C=2*pi*7

OpenStudy (anonymous):

so it would be 43.96 right ?

myininaya (myininaya):

your question says exact what you have is an apprioximation